OpenStudy (anonymous):

The line x-6y+11=0 bisects the joint of (8,-1)and (o,k).Find the value of k.

OpenStudy (anonymous):

find the midpoint of (8,-1) and (0,k) that is 4,(k-1)/2 now put this value in the line x-6y+11=0 because those coordinates must satisfy the line n you can get k ;)

OpenStudy (anonymous):

\[4-6\frac{ (k-1) }{ 2 }+11=0\]

OpenStudy (shubhamsrg):

Hello again auntie!